A 90 kg skier at the top of a 400 m icy hill slides down, through a 20 m patch of snow (µk= 0.25) and hits a 20 m long jump ramp at a 45° incline.  

a) (10 points)  How far from the edge of the ramp does the skier land (m)?

b) (5 points) How high will he get of the ground (max height (m)) ?
Answer #1

Solution - 4oom lom 145 h = 2osin 45° h = 14.4 mwhoo6l = hmax bo 8.647 60-19) To Vg hmax q 2 le ho SlL = x trix 5019 - 10 totshuiswh.go=n 325 t-Z1 = 7 a = blihl- 760.19 - 27

Please give positive response it helps me alot...thank you